Find the surface area of a cylinder with a base diameter of 4ft and a height of 4ft .
Write your answer in terms of Pie, and be sure to include the correct unit.

Answer by edjones(8007)

C=pi*d
=4*pi
A=C*h
=16*pi sq. ft.
.
Ed